Given the financial statements for Jones Corporation and Smith Corporation JONES CORPORATION Current Assets Liabilities Cash Accounts receivable Inventory $ 25,800 Accounts payable S188,000 80,400 88,700 54,200 Bonds payable (long term) Long-Term Assets Stockholders Equity $578,000 153,900 Gross fixed assets Common stock $150,000 70,000 104,400 $592,800 Less: Accumulated Paid-in capital 424,100 Retained earnings $592,800 Net fixed assets* Total assets Total liabilities and equity Sales (on credit) Cost of goods sold Gross profit Selling and administrative $1,717,000 782,000 $935,000 283,000 nse Depreciation expense Operating profit Interest expense Earnings before taxes $ 598,700 10,000 ax nse Net income $495,500 Use net fixed assets in computing fixed asset turnover tIncludes $11,100 in lease payments. SMITH CORPORATION Current Assets Liabilities Cash Marketable securities Accounts receivable Inventory $ 36,100 Accounts payable S 84,700 251,000 10,800 Bonds payable (ong term) 77,300 75,300 Long-Term Assets Stockholders Equity 75,000 30,000 78,100 $518,800 Gross fixed assets $576,000 256,700 Common stock Less: Accumulated Paid-in capital Net fixed assets* Total assets 319,300 Retained earnings $518.800 Total liabilities and equity Use net fixed assets in computing fixed asset turnover SMITH CORPORATION Sales (on credit) Cost of goods sold Gross profit Selling and administrative expense Depreciation expense Operating profit Interest expense Earnings before taxes lax expense Net income $1,540,000 1,110,000 $ 430,000 229,000 51,700 $ 149,300 29,500 119,800 40,100 Includes $11,100 in lease payments a. Compute the following ratios. (Use a 360-day year. Do not round intermediate calculations. Input your profit margin, return on assets, return on equity, and debt to total assets answers as a percent rounded to 2 decimal places. Round all other answers to 2 decimal places.)Jones Corp. Smith Corp Profit margin Return on assets investments Return on equity Receivable turnover Average collection times times eriod Inventory turnover Fixed asset turnover Total asset turnover Current ratio Quick ratio Debt to total assets Times interest earned Fixed charge coverage days times times times times times days times times times times times times times times times
Answer #1
Formula Jones Corporation Smith Corporation
Profit Margin=Net Income/Sales
Net Income=(A) $                  4,95,500.00 $                                79,700.00
Sales=(B) $               17,17,000.00 $                          15,40,000.00
Profit Margin=(A)/(B) 28.86% 5.18%
Return of Assets=Net Income/Total Assets
Net Income=(C ) $                  4,95,500.00 $                                79,700.00
Total Assets=(D ) $                  5,92,800.00 $                            5,18,800.00
Return of Assets=(C )/(D) 83.59% 15.36%
Return on Equity
Net Income/Equity
Net Income=(A) $                  4,95,500.00 $                                79,700.00
Equity
Common Stock $                  1,50,000.00 $                                75,000.00
Paid in Capital $                     70,000.00 $                                30,000.00
Retained Earnings $                  1,04,400.00 $                                78,100.00
Total Equity=(B) $                  3,24,400.00 $                            1,83,100.00
Return on Equity=(A)/(B) 152.74% 43.53%
Receivable Turnover
Sales/Debtors
Sales=(A) $               17,17,000.00 $                          15,40,000.00
Debtors=(B) $                     88,700.00 $                                77,300.00
Receivable Turnover=(A)/(B) 19.36 19.92 Times
Average collection period
360/Receivable Turnover
Days=(A) 360 360
Receivable Turnover=(B) 19.36 19.92
Average collection period=(A)/(B) 18.60 18.07 Days
Inventory Turnover
Cost of goods sold/Inventory
Cost of goods sold=(A) $                  7,82,000.00 $                          11,10,000.00
Inventory=(B) $                     54,200.00 $                                75,300.00
Inventory Turnover=(A)/(B) 14.43 14.74 Times
Fixed Assets Turnover
Sales/Fixed Assets
Sales=(A) $               17,17,000.00 $                          15,40,000.00
Fixed Assets=(B) $                  4,24,100.00 $                            3,19,300.00
Fixed Assets Turnover=(A)/(B) 4.05 4.82 Times
Total Assets Turnover
Sales=(A) $               17,17,000.00 $                          15,40,000.00
Total Assets=(B) $                  5,92,800.00 $                            5,18,800.00
Total Assets Turnover=(A)/(B) 2.90 2.97 Times
Current Ratio
Current Assets/Current Liabilites
Current Assets
Cash $                     25,800.00 $                                36,100.00
Marketable Securities $                                10,800.00
Accounts Receivable $                     88,700.00 $                                77,300.00
Inventory $                     54,200.00 $                                75,300.00
Total Current Assets=(A) $                  1,68,700.00 $                            1,99,500.00
Current Liabities
Accounts Payable $                  1,88,000.00 $                                84,700.00
Total Current Liabilites=(B) $                  1,88,000.00 $                                84,700.00
Current Ratio=(A)/(B) 0.90 2.36
Quick Ratio
Quick Assets
Cash $                     25,800.00 $                                36,100.00
Marketable Securities $                                10,800.00
Accounts Receivable $                     88,700.00 $                                77,300.00
Total Quick Assets=(A) $                  1,14,500.00 $                            1,24,200.00
Current Liabities
Accounts Payable $                  1,88,000.00 $                                84,700.00
Total Current Liabilites=(B) $                  1,88,000.00 $                                84,700.00
Quick Ratio=(A)/(B) 0.61 1.47
Debt to total Assets
Debt/Total Assets
Debt
Accounts Payable $                  1,88,000.00 $                                84,700.00
Bonds Payable $                     80,400.00 $                            2,51,000.00
Total Debt=(A) $                  2,68,400.00 $                            3,35,700.00
Total Assets=(B) $                  5,92,800.00 $                            5,18,800.00
Debt /Total Assets=(A)/(B) 45.28% 64.71%
Times Interest Earned
Earnings before interest & tax/Interest
Earnings before interest & tax=(A) $                  5,98,700.00 $                            1,49,300.00
Interest=(B) $                     10,000.00 $                                29,500.00
Earnings before interst &tax/Interest=(A)/(B) 59.87 5.06 Times